# An Empirical Study of LLM-as-a-Judge for LLM Evaluation: Fine-tuned Judge Model is not a General Substitute for GPT-4

### v3

@@ 3.2.4 No Benefits from Prompt Engineering
We validate both close-sourced and fine-tuned judges with the two strategies [^5] The detailed prompts are presented in Appendix A.2.. As shown in Table 6, the close-sourced models are improved by a large margin through both prompt engineering strategies. Conversely, the fine-tuned judges hardly benefit from these strategies, sometimes even experiencing severe performance decline. Specifically, in the case of CoT prompting, despite we modified the prompts for JudgeLM and PandaLM to generate CoT firstly, both models adhered to their original output format and failed to produce CoT. While there exist more intricate prompting strategies such as ChatEval Chan et al. (2024) or Branch-Solve-Merge Saha et al. (2023), we posit they can neither bring benefit to the fine-tuned judges, as they have lost their general instruction-following ability and are constrained to a singular output pattern.
3.3 The Essence of Fine-tuned Judge: A Task-specific Classifier
@@
    Vicuna-generation: It formulates the evaluation task in a generation-style, and the prediction head reuses the pretrained language model head, and is trained akin to the process of language modeling, based on the 7B version of Vicuna Chiang et al. (2023);
@@
    Vicuna-classification: It formulates the evaluation task as classification or regression, and the prediction head is newly initialized as a linear projection layer, and is decoupled from the language modeling process;
@@
Model SALAD-Bench JudgeLM-test PandLM-test Auto-J-test Average
accuracy F1 accuracy F1 accuracy F1 agreement
GPT-4-1106 98.75 65.55 84.24 72.83 75.78 71.51 56.90 78.92
JudgeLM-7B 82.81 57.46 82.39 72.97 68.17 65.18 45.33 69.68
+CascadedEval 96.35 64.56 84.53 71.69 75.08 71.19 54.09 77.51
Auto-J-13B 72.76 52.72 77.79 62.64 72.17 64.10 53.59 69.08
+CascadedEval 93.85 62.73 81.07 64.91 76.27 72.09 57.69 77.22
Table 8: Experiment results of our proposed CascadedEval. We use SE-Cali to quantify the confidence of fine-tuned judge models, and then allocate the less confident 50% to be evaluated by GPT-4.
Model Indicator SALAD-Bench JudgeLM-test PandLM-test Auto-J-test Average
accuracy F1 accuracy F1 accuracy F1 agreement
JudgeLM-7B random 82.50 57.36 80.40 71.87 67.54 65.20 43.25 59.42
perplexity 88.09 59.57 85.48 74.38 72.34 63.25 49.57 61.69
SE 95.10 63.62 89.40 78.09 77.96 64.01 55.89 65.40
SE-Cali 91.56 61.47 88.70 76.30 79.76 67.05 56.32 65.29
Auto-J-13B random 72.29 48.10 77.27 60.7 71.54 63.51 46.12 54.61
perplexity 77.47 52.56 80.28 61.85 75.55 65.84 50.35 57.65
SE 81.98 54.68 79.13 63.58 80.16 66.61 53.01 59.47
SE-Cali 82.08 54.75 80.19 63.58 81.36 70.00 52.30 60.16
Table 9: Comparison of different confidence indicators for the judge models. We split the test sets into halves based on different indicators, and report the performance of the judge on the half with higher scores.
Figure 4: The overall framework of CascadedEval. Conf. denotes the confidence score qualified by our indicator.
Notice for fine-tuning Vicuna-generation and Vicuna-classification models, we adopt the same prompt and hyper-parameters, with the only difference lying in the prediction method [^7] An illustration figure is presented in Appendix A.1..
As shown in Table 7, the classification model performs equally well as the generation model on both pairwise selection and pointwise grading. The formidable generative capabilities of LLMs hardly bring any improvement to the evaluation, as they are fitting to the same group of data. Moreover, the DeBERTa-based classifier achieves comparable performance with the LLM-based evaluators [^8] The only exception is on Auto-J-test, which is possibly due to a large proportion of the test data exceeds 512 (the maximum context length of DeBERTa)., which might be argued for that the encoder-only architecture is more suitable for classification.
@@
4 Integrate Fine-tuned Judge with GPT-4
4.1 CascadedEval
Despite the limitations revealed in our study, we aim not to disregard the significance of fine-tuned judges entirely. While LLMs are becoming increasingly prevalent, task-specific models are still wildly used in real applications, considering their superior in-domain performance and cost-effectiveness. Therefore, to make the most of fine-tuned judges for LLM evaluation, we propose a novel method, CascadedEval, integrating proprietary models such as GPT-4 to compensate for the limitations.
The framework of CascadedEval is shown in Figure 4. When applying fine-tuned judge for evaluation, CascadedEval derive the confidence scores as a by-product and allocate the less confident samples to be re-evaluated by GPT-4.
The key of our method is the confidence indicator. While previous works resort to perplexity or self-reflection to quantify the confidence Jung et al. (2024), in this work, we propose an effective confidence indicator based on softmax probability distribution. Given an instruction $x$ and a fine-tuned judge model with parameters $\theta$, the confidence of evaluation $y$ can be estimated as:
$\textrm{SE}(y|x,\theta)=-\frac{1}{T}\sum_{t=1}^{T}\sum_{v=1}^{V}p_{\theta}(y_{t}^{v})\mathrm{log}p_{\theta}(y_{t}^{v}),$ (1)
where $p(y_{t})$ represents the conditional distribution $p(y_{t}|x,y_{<t},\theta)$ at each decoding step, $T$ is the response length, and $V$ is the vocabulary size. If the majority of the probability mass is concentrated on a limited number of vocabulary words, it indicates that the model is confident and the evaluation is more likely to be accurate, and vice versa.
As we would like to quantify whether the sample lies in the task-specific fine-tuning scope, we further calibrate the SE as follows:
[…]
where $\theta_{j}$ denotes the fine-tuned judge model, and $\theta_{b}$ denotes its corresponding foundation model. By calibration, we aim to exclude the influence of foundation model, thus modeling solely the confidence instilled by the task-specific fine-tuning process.
With the effective confidence indicator, CascadedEval enhances the accuracy of fine-tuned judges by routing the samples that fall outside the fine-tuning scope to GPT-4. Conversely, compared to using GPT-4 for all evaluations, CascadedEval significantly reduces the API expense without leading to performance degradation. This is combining the best of both worlds.
4.2 Experiments
In this part, we verify the effectiveness of our proposed CascadedEval on both in-domain and out-of-domain meta-evaluation testsets. As shown in Table 8, by allocating 50% of the less confident samples to GPT-4, CascadedEval achieves an accuracy on par with GPT-4, even surpassing GPT-4 sometimes. This verifies that the fine-tuned judges are highly effective when the evaluation sample aligns with the fine-tuning scope. With our proposed confidence indicator, the degree of this alignment can be accurately quantified, allowing for the targeted processing of less confident cases using enhanced proprietary models.
We also conducted an ablation study, where we control the proportion of samples allocated to GPT-4. As shown in Figure 5, with more GPT-4 involvement, the evaluation accuracy firstly increases and then slightly decreases. This notifies that there exists a balance in CascadedEval between fine-tuned judge and GPT-4, which should be carefully monitored to ensure maximum benefit.
Moreover, to verify the effectiveness of the confidence indicator, we re-conduct the cross-validation in Section 3, wherein the test set was split into halves based on different indicators [^9] We did not compare with self-reflection, as the judge model has failed the general instruction-following ability.. As shown in Table 9, our proposed indicators managed to select the samples with higher accuracy. This verifies the effectiveness of softmax distribution-based indicator for confidence estimation.
In an additional experiment, we group the test set into five distinct buckets based on the SE-Cali scores of each sample, and evaluate the accuracy for each bucket separately. As shown in Figure 6, the accuracy on each bucket exhibit a strong correlation with the SE-Cali score. This underscores the efficacy of SE-Cali as a reliable confidence indicator of the judge model.
Figure 5: The variation of accuracy on PandaLM-test with more samples allocated to GPT-4.
Figure 6: Accuracy of Auto-J on SALAD-Bench when applied to different buckets of data grouped by SE-Cali scores, highest in bucket 1 while lowest in bucket 5.
@@ 5 Conclusion
In this work, we conduct an empirical study on the judge models for LLM evaluation. As revealed in our experiments, despite achieving superior evaluation performance on the in-domain testset, the fine-tuned judge models still present several limitations compared with GPT-4.
As the potential of LLM extends beyond boundaries, there will always be new domains and tasks not covered by the fine-tuning process. Therefore, to draw a conclusion, the fine-tuned judge model cannot serve as a general substitution for GPT-4 in terms of LLM evaluation. It is advisable to exercise caution when leveraging fine-tuned judge models for evaluation in real applications, and introduce proprietary models when necessary.
@@
Our work still has some limitations: 1) The CascadedEval proposed in our paper can be combined with human annotator in a semi-supervised manner, to further improve the LLM evaluation pipeline. For example, we can select the least confident samples to be evaluated by human evaluator, therefore ensuring the evaluation accuracy with minimal expense. Active learning can also be introduced, leading to more targeted fine-tuning on the less confident samples. Due to time and resource constraints, we have to leave this as future work. 2) The work of Zeng et al. (2023) is only a general assessment of evaluator bias, and we did not include fine-grained assessment for different biases, such as position bias Wang et al. (2023a), verbosity bias Saito et al. (2023), etc. 3) Due to time constraints, we did not incorporate manual inspection into the meta-evaluation process. Including human evaluators would enhance the credibility of our claims.
@@ A.1 Training Settings
As mentioned in Section 3, we fine-tune our judge models based on the four groups of data (JudgeLM Zhu et al. (2023b), PandaLM Wang et al. (2024), Auto-J Li et al. (2023a), Prometheus Kim et al. (2023)), both in generation-style and in classification-style, for the purpose of comparison.
A.2 Prompt Templates
As mentioned in Section 3, we take the publicly released checkpoints of the four fine-tuned judge models and validate their performance. To make a fair comparison, we make minimal modifications to their pre-defined prompts, to adapt them to different scenarios. The specific prompts designed for different sections are listed as follows:
  1. 1.
    For Section 3.2.1, we adopt the prompts presented in Figure 8 to 15 for cross validation. Notice for JudgeLM and PandaLM, their predefined prompts are in the form of pairwise selection, and we make slight modifications to apply them on pointwise grading. For Prometheus, the predefined prompt is in the form of pointwise grading, and we make slight modifications to apply it on pairwise selection. For Auto-J, they predefined prompts both for pairwise selection and pointwise grading. We also adopt the prompts presented from Figure 16 to 19 on MT-Bench, which are all adapted to multi-turn evaluation.
  2. 2.
    For Section 3.2.2, we adopt the prompts presented in Figure 8, 10, 12 and 14, as LLMBar is pair-wise selection.
  3. 3.
    For Section 3.2.3, we adopt the prompts presented in Figure 20 to 23 for JudgeLM, PandaLM and Auto-J, respectively. For Prometheus, as its original prompt comprises of scoring rubrics, we simply define the corresponding rubrics for different benchmarks.
  4. 4.
    For Section 3.2.4, we adopt the prompts presented in Figure 24 and Figure 25 for chain-of-thought prompting.

### v4

@@ 2.4 Can not Benefit from CoT and ICL
As shown in Table 7, while the close-sourced proprietary models are improved by a large margin through both prompt engineering strategies, the fine-tuned judges hardly benefit from these strategies, sometimes even experiencing severe performance decline. Specifically, in the case of CoT prompting, despite we modified the prompts for JudgeLM and PandaLM to generate CoT firstly, both models failed to produce CoT and adhered to their original output format, as they have lost their general instruction-following ability.
We also evaluated the impact of different CoT sources based on JudgeLM-7B. We first swapped the positions of scores and CoT in the training data, and then fine-tuned the base model with or without CoT using the same hyperparameters. Additionally, we utilized o1-preview-0912222platform.openai.com/docs/models/o1 to generate CoT for the original scores through hint-driven prompting Srivastava et al. (2023), and subsequently fine-tuned the model with this annotated CoT.
As demonstrated in Table 6, fine-tuning with either the original CoT or the o1-generated CoT resulted in a degradation of model performance compared to the model fine-tuned without CoT. Notably, the o1-generated CoT led to a more severe performance drop. This clearly indicates that even high-quality CoT did not introduce any improvement to the fine-tuned judge.
3 The Essence of Fine-tuned Judge: A Task-specific Classifier
@@
    Vicuna-generation Chiang et al. (2023): It formulates the evaluation task in a generation-style, and the prediction head reuses the pretrained language model head;
@@
    Vicuna-classification: It formulates the evaluation task as classification or regression, and the prediction head is newly initialized as a linear projection layer;
@@
As shown in Table 5, the classification model performs equally well as the generation model. The formidable generative capabilities of LLMs hardly bring any improvement to the evaluation, as they are fitting to the same group of data. Moreover, the DeBERTa-based classifier achieves comparable performance with the LLM-based evaluators [^4] The only exception is on Auto-J-test, which is possibly due to a large proportion of the test data exceeds 512., which might be argued for that the encoder-only architecture is more suitable for classification.
@@ 4 Conclusion
Although the fine-tuned models demonstrate superior performance on in-domain test sets, they still have several limitations compared to GPT-4. While increasing the fine-tuning data could possibly mitigate some of the limitations, as the potential of LLM extends beyond boundaries, there will always be new domains and tasks that are not covered by the fine-tuning scope. Therefore, the fine-tuned judge model cannot replace GPT-4 as a universal evaluator for LLMs, and should be used judiciously by watching the domain and task adaptability.
@@
Our work still has some limitations: 1) Due to time limitation, we did not present a possible solution to mitigate the limitations of fine-tuned judge models. We will investigate related method in the future. 2) The work of Zeng et al. (2023) is only a general assessment of evaluator bias, and we did not include fine-grained assessment for different biases, such as position bias Wang et al. (2023a), verbosity bias Saito et al. (2023), etc. 3) Due to time constraints, we did not incorporate manual inspection into the meta-evaluation process. Including human evaluators would enhance the credibility of our claims.
Acknowledgements
This work is supported by National Natural Science Foundation of China (62276077, 62376075, 62376076).
@@ A.1 Training Settings
As mentioned in Section 1, we fine-tune our judge models based on the four groups of data (JudgeLM Zhu et al. (2024), PandaLM Wang et al. (2024), Auto-J Li et al. (2024a), Prometheus Kim et al. (2024)), both in generation-style and in classification-style, for the purpose of comparison.
@@
A.2 Prompt Templates
As mentioned in Section 1, we take the publicly released checkpoints of the four fine-tuned judge models and validate their performance. To make a fair comparison, we make minimal modifications to their pre-defined prompts, to adapt them to different scenarios, as listed as follows.
For Section 2.1, we adopt the prompts presented in Figure 5 to 12 for cross validation. Notice for JudgeLM and PandaLM, their predefined prompts are in the form of pairwise selection, and we make slight modifications to apply them on pointwise grading. For Prometheus, the predefined prompt is in the form of pointwise grading, and we make slight modifications to apply it on pairwise selection. For Auto-J, they predefined prompts both for pairwise selection and pointwise grading. We also adopt the prompts presented from Figure 13 to 16 on MT-Bench, which are all adapted to multi-turn evaluation. We adopt the prompts presented in Figure 21 and Figure 22 for chain-of-thought prompting.
For Section 2.2, we adopt the prompts presented in Figure 5, 7, 9 and 11, as LLMBar is in the form of pair-wise selection.
For Section 2.3, we adopt the prompts presented in Figure 17 to 20 for JudgeLM, PandaLM and Auto-J, respectively. For Prometheus, as its original prompt comprises of scoring rubrics, we simply define the corresponding rubrics for different benchmarks. As HaluEval and ToxicChat are both binary classifications, we apply Auto-J and Prometheus with pointwise grading and conduct a grid search to determine the classification threshold. On the other hand, as SALAD-Bench is a pairwise classification, we apply pairwise selection models, namely JudgeLM, PandaLM, and Auto-J to select a better response.